OpenStudy (anonymous):

If there are 22 marbles and you pick out 2, what are the odds you'll get 2 red marbles? 4-red 6-white 5-blue 2-white 3-purple

OpenStudy (anonymous):

4x + 6y = 14 -2x + 2y = -2

OpenStudy (anonymous):

Idk..

OpenStudy (anonymous):

so there is 22 marbles and then you would take the 4 red marbles and it would be 4/22 but can be simplified to 2/11

OpenStudy (texaschic101):

wouldn't that be for just 1 red marble ?

OpenStudy (anonymous):

that only equals 20 your missing two somewhere

OpenStudy (anonymous):

yeah but i was waiting for him to reply to finish explaing @texaschic101 and @TessaMarie08181997 is right they only add up to 20 so somewhere you are missing 2

OpenStudy (anonymous):

Whoops I meant 5 purple!

OpenStudy (texaschic101):

4/22 = 2/11....that would be for 1 red marble...right ?

OpenStudy (anonymous):

yes but you will get the same thing tho

OpenStudy (anonymous):

its 4/22 = 1/11 and in decimal its 0.181818... repeating

OpenStudy (anonymous):

@josie_2015 isnt that bright of a light ya'll..

OpenStudy (texaschic101):

@TessaMarie08181997 ....how does 4/22 = 1/11 ?

OpenStudy (anonymous):

i ment 2/11 sorry i was writing one on my math worksheeet

OpenStudy (texaschic101):

the first marble is 2/11.....is the second marble would be 4/21...right ? Or am I way off ?

OpenStudy (anonymous):

he starts out with 4/22 when he takes two out it will be 2/20

OpenStudy (texaschic101):

oh.....ok :)

OpenStudy (anonymous):

yeah